For a compound circuit R series with XL , R = 3 omega and XL = 6O omega connected with parallel branches R = 2 omega and XC = 8 omega E=30V angle 0^ Find source (IT) .


1
Expert's answer
2021-04-08T07:33:11-0400

"I=\frac{E}{Z_1+Z_2}"


"Z_1=\sqrt{R_1^2+X^2_{L1}}=\sqrt{3^2+60^2}=60.07\ \Omega"


"\frac{1}{Z_2}=\sqrt{\frac{1}{R_2^2}+\frac{1}{X_{C2}^2}}=\sqrt{\frac{1}{2^2}+\frac{1}{8^2}}=\frac{\sqrt{17}}{8} \ 1/\Omega"

"Z_2=8/\sqrt{17}=1.94\ \Omega"


"I=\frac{30}{60.07+1.94}=0.48\ A"